OpenStudy (anonymous):

200 raffle tickets are sold at $3 each. 1 prize of $100 is to be awarded. Raul purchases 1 ticket. (a)determine his expected value...(b) determine the fair price of the ticket...@mathslover

OpenStudy (anonymous):

The probability he wins is \(1/200\). The probability he loses is \(199/200\).

OpenStudy (anonymous):

If he wins, he nets \( $100-$3\). If he loses, he nets \(-$3\).

OpenStudy (anonymous):

Expected value is a weighted average of the net value of each outcome weighted by the probability of said outcome.

OpenStudy (anonymous):

ok...so how do you determine the fair price

OpenStudy (anonymous):

Fair price would mean the expected value is zero.

OpenStudy (anonymous):

so replace \(3\) with \(x\), set the expression equal to \(0\), and solve for \(x\).

OpenStudy (anonymous):

You want: \[ E[X] = 0 \implies \sum x_i\Pr(X=x_i)=0 \]

OpenStudy (anonymous):

Now, we've established that: \[ E[X] = \frac{$100-$3}{200} + \frac{199(-$3)}{200} \]If we let the price be \(p\), then: \[ E[X] = \frac{$100-p}{200} + \frac{199(-p)}{200} \]We want \(E[X] =0 \), so: \[ \frac{$100-p}{200} + \frac{199(-p)}{200}=0 \]

OpenStudy (anonymous):

When the expected value is \(0\), it means that neither the raffle sponsors, nor the players, profits from the raffle.

OpenStudy (anonymous):

can u help me with another one?

OpenStudy (anonymous):

Open a new question.